Circular No. 8437

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004fz IN NGC 783 Independent discoveries of an apparent supernova in NGC 783 have

been reported by T. Boles, Coddenham, England, and by R. Arbour, South Wonston, England, from unfiltered CCD images. Boles provides the following position from his image taken on Nov 14.085 UT with a 0.35-m reflector: α = 2^{h}01^{m}06^{s}.41, δ = +31^{o}52′44′′.2, which is ≈ 2′′.4 west and 13′′.0 south of the center of NGC 783; Arbour provides position end figures 06^{s}.41, 44′′.3. Available magnitudes for SN 2004fz: 1991 Sept. 16, [21.0 (Digitized Sky Survey red plate, via Boles); 1993 Sept. 14, [21.5 (Digitized Sky Survey blue plate, via Boles); Sept. 8, [19.5 (Boles); Oct. 5, [19.5 (Boles); Nov. 14.084, 15.9 (Arbour); 14.085, 16.5 (Boles); 14.744, 16.0 (Arbour). Arbour also provides limiting magnitudes from Palomar Sky Survey plates: 1993.7, blue, [19.7; 1994.4, red, [20.1.

SUPERNOVAE 2004fa, 2004fd, 2004fg, 2004fw, 2004fy M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrogram (range 340–710 nm) of SN 2004fd (cf. IAUC 8423), obtained by Modjaz on Nov. 13.33 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a supernova of type Ia, with a spectral-feature age of about 20 days after B maximum. Adopting the NED recession velocity of 5190 km/s for the host galaxy, the supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm), is 9000 km/s.

A. V. Filippenko and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 320–940 nm), obtained on Nov. 14 UT with the 10-m Keck I telescope (+ LRIS), shows that SN 2004fa (IAUC 8420), SN 2004fd, SN 2004fg (IAUC 8425), SN 2004fw (IAUC 8431), and SN 2004fy (IAUC 8435) are of type Ia, with respective spectral-feature ages (Riess et al. 1997, A.J. 114, 722) of 2 months, 19 days, 0 day, 8 days, and 11 days past maximum brightness (with an uncertainty of ± 2 days at early times). The expansion velocities of the ejecta, as measured from the minimum in the Si II 635.5-nm absorption trough, is ∼ −11000, −7700, and −9500 km/s for SN 2004fg, SN 2004fw, and SN 2004fy, respectively. Interstellar Na I D absorption at the redshifts of the host galaxies is visible in the spectra of SN 2004fg and SN 2004fw (0.1-nm equivalent width in both cases).

2004 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green